OpenStudy (anonymous):

Use the substitution method to solve the following system of equations. 4x - y = 3 7x - 9y = -2 A) (1, 1) B) (6, -3) C) (6, 1) D) (1, -3)

OpenStudy (mathstudent55):

Do you know how the substitution method works?

OpenStudy (anonymous):

from eq 1..find out value of y..what you get..??write here...

OpenStudy (anonymous):

im not sure /: to either of your questions sorry @mathstudent55 @zaibali.qasmi

OpenStudy (anonymous):

i got (1, -3) for final answer... I'm not sure it is right

OpenStudy (anonymous):

@chariety101 from eq 1 find the value of y and put it in then eq 2..see below how is tis..

OpenStudy (anonymous):

its answer is (1,1)

OpenStudy (anonymous):

i must have made a mistake somewhere

OpenStudy (anonymous):

4x-y=3=>y=4x-3 now put it in eq 2 so it will be.. 7x-9(4x-3)=-2 =>7x-36x+27=-2 =>-29x=-29 =>x=1 now x=1 put in y=4x-3 so y=4*1-3=1 so (x,y)=(1,1)

OpenStudy (anonymous):

yuuupp..see the sol below..

OpenStudy (anonymous):

get it..?

OpenStudy (anonymous):

yes yes i get it thank you!!

OpenStudy (anonymous):

its oky dear...

OpenStudy (mathstudent55):

1. The method of substitution is to first solve one equation for one variable. 2. Then insert what that variable is equal to into that variable in the other equation. 3. Then solve the equation for the variable. 4. Then substitute the known variable into one of the equations and find the other variable.

OpenStudy (mathstudent55):

4x - y = 3 7x - 9y = -2 1. Solve the first equation for y. 4x - y = 3 Subtract 4x from both sides: -y = -4x + 3 Divide both sides by -1 y = 4x - 3 2. Now insert what y is equal to into y of the other equation. 7x - 9y = -2 7x - 9(4x - 3) = -2 3. Now solve the above equation for x 7x - 36x + 27 = -2 -29x = -29 x = 1 4. Now insert 1 for x in either of the original equations and solve for y. I'll use the first equation. 4x - y = 3 4(1) - y = 3 4 - y = 3 -y = -1 y = 1 Answer is x = 1 and y = 1, or (1, 1)